OpenStudy (anonymous):

One number is 4 less than a second number. Twice the second number is 34 more than 4 times the first. Find the two numbers. A) -12 and -8 B) 9 and 13 C) -14 and -10 D) -13 and -9

OpenStudy (anonymous):

x-y=4 4x-y=34 is this right

OpenStudy (anonymous):

the first part is right

OpenStudy (anonymous):

thje second is 2y-4x=34

OpenStudy (anonymous):

oh Thank you!

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

wait i think its x-y= -4

OpenStudy (anonymous):

depends on which you are using as second number

OpenStudy (anonymous):

im so lost! Itried plugging in the options none of them worked

OpenStudy (anonymous):

if your second number is y than y-x=4, than x-y=-4

OpenStudy (anonymous):

because the x is 4 less than the y

OpenStudy (anonymous):

oh ok 2(9)-4(13)= -34

OpenStudy (anonymous):

9-13=-4

OpenStudy (anonymous):

so would b be the correct answer!

OpenStudy (anonymous):

one sec let me test it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yea it works one sec ill show you

OpenStudy (anonymous):

ok

OpenStudy (anonymous):